# Structure-preserving approximation of the non-isothermal Cahn-Hilliard system

### v1

@@ Abstract
We propose and analyze a structure-preserving approximation of the non-isothermal Cahn-Hilliard equation using conforming finite elements for the spatial discretization and a problem-specific mixed explicit-implicit approach for the temporal discretization. To ensure the preservation of structural properties, i.e. conservation of mass and internal energy as well as entropy production, we introduce a suitable variational formulation for the continuous problem, based on the entropy equation. Analytical findings are supported by numerical tests, including convergence analysis.
@@ 1 Introduction
The Cahn-Hilliard equation is well-known in physics to describe conservative phase transitions like magnetism [21] and spinodal decomposition [10, 5], particularly in binary systems like alloys or liquid mixtures. These phenomena typically involve the separation of two phases when the system is quenched below a critical temperature. In many applications, the system is forced into a critical regime, often under the assumption of constant temperature, i.e., an isothermal system. The impact of non-isothermal conditions in such processes have been considered in [27, 34, 35]. Cahn-Hilliard models have also been applied to tumor growth, as seen in isothermal models [17] and their non-isothermal counterparts [25]. Additionally, non-isothermal phase-field models are of interest in material science, particularly in non-isothermal sintering and powder bed fusion [41, 40, 31]. In these scenarios, the system can cross the critical temperature barrier locally in both directions, necessitating the study of non-isothermal variations of the Cahn-Hilliard equation. In this work we consider the non-isothermal Cahn-Hilliard equation (NCH) in the spirit of Pawlow and Alt [2] given by
$\partial_{t}\phi-\operatorname{div}(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta})$ $=0,$ (1)
@@
$\partial_{t}e(\phi,\vartheta)-\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})$ $=0.$ (3)
Here $\phi$ is the phase-field variable distinguishing between phase A, where $\phi=0$ and phase B, where $\phi=1$, $\tfrac{\mu}{\vartheta}$ is the chemical potential and $\vartheta$ is the temperature. The matrices $\mathbf{M},\mathbf{K},\mathbf{C}$ are the mobility matrix, heat conductivity matrix and a cross-coupling matrix. The above system is complemented with initial data and stated on a periodic space domain $\Omega$ and the finite time interval $(0,T)$. In addition the system is subjected to the second law of thermodynamics, i.e. entropy is non-decreasing in time. The connection of the chemical potential $\frac{\mu}{\vartheta}$, the internal energy density $e$ and the entropy density $s$ is given a Helmholtz free energy of the following form
$F(\phi,\vartheta):=\frac{\gamma\vartheta}{2}|\nabla\phi|^{2}+f(\phi,\vartheta).$ (4)
This allows us to show that the system satisfies the following structural properties dictated by thermodynamics.
@@
$\int_{\Omega}s(\phi(t),\vartheta(t))=\int_{\Omega}s(\phi(0),\vartheta(0))+\int_{0}^{t}\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})$
with the entropy production
@@
$\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})=\int_{\Omega}\begin{pmatrix}\nabla\frac{1}{\vartheta}&\nabla\frac{\mu}{\vartheta}\end{pmatrix}\begin{pmatrix}\mathbf{K}&-\mathbf{C}\\ -\mathbf{C}&\mathbf{M}\end{pmatrix}\begin{pmatrix}\nabla\frac{1}{\vartheta}\\ \nabla\frac{\mu}{\vartheta}\end{pmatrix}.$
Under usual compatibility assumptions of the Onsager matrix
$\mathbf{L}=\begin{pmatrix}\mathbf{K}&-\mathbf{C}\\ -\mathbf{C}&\mathbf{M}\end{pmatrix},$
which is assumed to be positive (semi)-definite, the total entropy is produced as time evolves.
In the literature there have been several different generalizations of phase-field models, non-conserved or conserved, for the non-isothermal setting; see [32, 24] for some general ansatz and a quite intensive overview. The well-known generalization for non-conserved phase-field models was introduced by Penrose and Five [33], while for a conserved phase-field model by Pawlow and Alt [2]. In the latter the system driving force is the entropy. Both models and related systems can also be derived within the well-known GENERIC or SNET formalism, see [18]. Modeling approaches based on the micro-force balance in the spirit of Gurtin [23] can be found in [30, 29] and the least action principle in [14]. Caginalp derived models with finite speed propagation of temperature, see [9, 13]. We may also note that these systems can be supplemented with (quasi)-incompressible flow, see [1, 15]. Colli et al. have studied several models from an analytical point of view, see [4, 12, 13]. Existence of weak solutions and further analysis for the model of Pawlow and Alt can be found in [3, 26]. Note that these results are the only available theoretical results dealing with the cross-coupling terms.
@@
To numerically approximate such models, a purely implicit Euler method was proposed by [32], using the inverse temperature as the state variable. This approach was generalized to an energy-stable method by González et al. [20] using the average vector field methods [19]. Recently, Brunk et al. [8, 7] introduced structure-preserving approximations based on this transformation. The system under consideration possesses a gradient structure. For the Cahn-Hilliard and Allen-Cahn equations, common numerical methods include convex-concave splitting [38], the scalar auxiliary variable (SAV) approach [37, 28], and Energy Quadratisation (EQ) [11, 42]. The SAV and EQ methods introduce auxiliary variables that are either time-dependent or space-time-dependent, thereby relaxing the driving functional energy or entropy. In the case of the internal energy equation, the Energy Quadratisation is used in [22, 39].
@@
The aim of this work is to study the variational properties of the NCH system (1)-(3) and its systematic discretisation in space and time using the standard temperature as state variable. The novelties of this manuscript are
@@
  • Problem-adapted reformulation, choosing the entropy equation instead of the internal energy equation, to reveal the variational structure.
  • Structure-preserving approximation via suitable application of standard discretisations in space and time.
@@
The manuscript is structured as follows. In Section 2, we introduce relevant notation and reformulate the system by exchanging the internal energy equation by its equivalent entropy equation. Afterwards, in Section 3, we will use the reformulated system to derive a variational formulation which is suitable for a direct approximation by a standard finite element method. In Sections 4 and 5, the fully discrete scheme using mixed explicit-implicit time discretisation and a convex-concave splitting of the driving potential is proposed and analysed. Finally, we present a convergence test and illustrate the evolution of the system through an application-inspired example in Section 6.
@@ 2 Preliminaries
Before we present our discretization method and main results in detail, let us briefly introduce our notation and main assumptions, and recall some basic facts.
@@ 2.1 Notation
The system (1)–(3) is investigated on a finite time interval $(0,T)$. To avoid the discussion of boundary conditions, we consider a spatially periodic setting, i.e.,
@@
    $\Omega\subset\mathbb{R}^{d}$, $d=1,2,3$ is a cube and identified with the $d$-dimensional torus $\mathcal{T}^{d}$, i.e. periodic boundary conditions are applied. Moreover, functions on $\Omega$ are assumed to be periodic throughout the paper.
[…]
As usual, we denote by $L^{p}(a,b;X)$, $W^{k,p}(a,b;X)$, and $H^{k}(a,b;X)$, the Bochner spaces of integrable or differentiable functions on the time interval $(a,b)$ with values in some Banach space $X$. If $(a,b)=(0,T)$, we omit reference to the time interval and briefly write $L^{p}(X)$. The corresponding norms are denoted, e.g., by $\|\cdot\|_{L^{p}(X)}$ or $\|\cdot\|_{H^{k}(X)}$.
For later use we will introduce the following theorem from [16]:
@@
Let N and M be two positive integers and V be defined as follows:
$V=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ y>0\},$
where the notation $\mathbf{y}>c$ means that each component of $\mathbf{y}$ is greater than $c$, $c\in\mathbb{R}$. Let $\mathbf{b}\in\mathbb{R}^{N}\times\mathbb{R}^{M}$ and $\mathbf{g}$ and $\mathbf{G}$ be two continuous functions respectively from $V$ and $V\times[0,1]$ to $\mathbb{R}^{N}\times\mathbb{R}^{M}$ satisfying:
@@
    $\mathbf{G}(\cdot,1)=\mathbf{g}(\cdot);$
@@
    $\forall\alpha\in[0,1]$, if $\mathbf{v}$ is such that $\mathbf{G}(\mathbf{v},\alpha)=\mathbf{b}$ then $\mathbf{v}\in W$ , where $W$ is defined as follows:
    $W=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M}\leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ \|\mathbf{x}\|<C_{1}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \varepsilon<\mathbf{y}<C_{2}\}$
[…]
@@
    the topological degree of $\mathbf{G}(\cdot,0)$ with respect to $\mathbf{b}$ and $W$ is equal to $d_{0}\neq 0$.
Then the topological degree of $\mathbf{G}(\cdot,1)$ with respect to $\mathbf{b}$ and $W$ is also equal to $d_{0}\neq 0$. Consequently, there exists at least one solution $\mathbf{v}\in W$ such that $\mathbf{g}(\mathbf{v})=\mathbf{b}$.
@@ 2.2 Assumptions
In this subsection we collect the minimal assumptions for the rest of the manuscript.
@@
    The interface parameter $\gamma\in\mathbb{R}$ is positive.
@@
    The Onsager matrix $\mathbf{L}(\phi,\nabla\phi,\vartheta)=:\mathbf{L}\in\mathbb{R}^{2d\times 2d}$ is assumed to be positive definite. It means that for positive constants $\lambda_{0},\lambda_{1}>0$ we have
    $\lambda_{1}|\xi|^{2}\geq\mathbf{\xi}^{\top}\mathbf{L}\mathbf{\xi}\geq\lambda_{0}|\xi|^{2},\quad\forall\mathbf{\xi}\in\mathbb{R}^{2d}.$
@@
    For a smooth driving potential $f(\cdot,\cdot):\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ we assume that for every fixed $\phi$ the potential $f(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is concave and goes to infinity as $\vartheta\to 0$. Furthermore, for every fixed ϑ+italic-$\vartheta\in\mathbb{R}_{+}$ the potential $f(\cdot,\vartheta)$ can be decomposed in a convex and a concave part. This separation is denoted by $f_{\text{vex}},f_{\text{cav}}.$
@@ 2.3 Structural identities
$\langle\partial_{t}\phi,1\rangle=\langle\partial_{t}e(\phi,\vartheta),1\rangle=0,\quad\langle\partial_{t}s(\phi,\vartheta),1\rangle=\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})\geq 0.$ (5)
@@
$\partial_{t}s(\phi,\vartheta)$ $=\partial_{t}\tfrac{e(\phi,\vartheta)-F(\phi,\vartheta))}{\vartheta}=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{t}F(\phi,\vartheta)-\tfrac{e(\phi,\vartheta)-F(\phi,\vartheta)}{\vartheta^{2}}\partial_{t}\vartheta$
$=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{\phi}F(\phi,\vartheta)\partial_{t}\phi-\tfrac{1}{\vartheta}\partial_{\vartheta}F(\phi,\vartheta)\partial_{t}\vartheta-\tfrac{s(\phi,\vartheta)}{\vartheta}\partial_{t}\vartheta.$
Using the relation $s(\phi,\vartheta)=-\partial_{\vartheta}F(\phi,\vartheta)$ and substituting in (4) leads to
$\partial_{t}s(\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{\phi}f(\phi,\vartheta)\partial_{t}\phi-\gamma\nabla\phi\cdot\nabla\partial_{t}\phi.$
@@
$\partial_{t}s(\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})-\tfrac{\mu}{\vartheta}\partial_{t}\phi-\gamma\operatorname{div}(\partial_{t}\phi\nabla\phi).$
@@
$\partial_{t}s(\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})-\tfrac{\mu}{\vartheta}\operatorname{div}(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta})$
@@
$\partial_{t}s(\phi,\vartheta)$ $=\nabla\tfrac{1}{\vartheta}\cdot\mathbf{K}\nabla\tfrac{1}{\vartheta}-2\nabla\tfrac{1}{\vartheta}\cdot\mathbf{C}\nabla\tfrac{\mu}{\vartheta}+\nabla\tfrac{\mu}{\vartheta}\cdot\mathbf{M}\nabla\tfrac{\mu}{\vartheta}$
$\quad-\operatorname{div}\left(\gamma\partial_{t}\phi\nabla\phi+\tfrac{1}{\vartheta}\mathbf{K}\nabla\tfrac{1}{\vartheta}-\tfrac{1}{\vartheta}\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\tfrac{\mu}{\vartheta}\mathbf{C}\nabla\tfrac{1}{\vartheta}+\tfrac{\mu}{\vartheta}\mathbf{M}\nabla\tfrac{\mu}{\vartheta}\right).$ (6)

### v2

@@ Abstract
We present and investigate a structure-preserving approximation of the non-isothermal Cahn-Hilliard equation, employing conforming finite elements for spatial discretisation and a tailored mixed explicit-implicit scheme for time integration. To guarantee the preservation of key structural properties, namely mass and internal energy conservation, along with entropy production, we formulate the continuous problem within an appropriate variational framework based on the entropy equation. Our analytical results are validated through numerical experiments, including a convergence study
@@ 1 Introduction
The Cahn-Hilliard equation is a well-established model in physics for describing conservative phase transitions, such as magnetism [21] and spinodal decomposition [10, 5], particularly in binary systems like alloys or liquid mixtures. These transitions typically occur when a system is quenched below a critical temperature, prompting phase separation. In many applications, the system is studied under isothermal conditions, assuming a constant temperature. However, the influence of nonisothermal effects has gained attention in recent work [27, 35, 34]. Cahn-Hilliard-type models have also been applied beyond materials science, including tumour growth modelling, both in isothermal [17] and nonisothermal contexts [25]. In materials science, non-isothermal phase-field models are particularly relevant for processes such as sintering and powder bed fusion [41, 40, 31], where local temperature fluctuations can drive the system above or below critical thresholds. These scenarios necessitate non-isothermal variants of the Cahn-Hilliard equation. In this work, we study the non-isothermal Cahn-Hilliard (NCH) equation in the framework of Pawlow and Alt [2], given by the system:
$\partial_{t}\phi-\operatorname{div}\left(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta}\right)$ $=0,$ (1)
@@
$\partial_{t}e(\phi,\vartheta)-\operatorname{div}\left(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta}\right)$ $=0.$ (3)
Here $\phi$ is the phase-field variable distinguishing between phase A ($\phi=0$) and phase B($\phi=1$), $\tfrac{\mu}{\vartheta}$ is the chemical potential and $\vartheta$ is the absolute temperature. The matrices $\mathbf{M}=\mathbf{M}(\phi,\nabla\phi,\vartheta),\mathbf{K}=\mathbf{K}(\phi,\nabla\phi,\vartheta)$ and $\mathbf{C}=\mathbf{C}(\phi,\nabla\phi,\vartheta)$ represent the mobility, thermal conductivity, and cross-coupling coefficients and may depend on $\phi,\nabla\phi$ and $\vartheta$. The system is posed on a periodic spatial domain $\Omega$ over a finite time interval $(0,T)$, supplemented by initial conditions and subject to the second law of thermodynamics, ensuring non-decreasing entropy over time. The Helmholtz free energy density $F(\phi,\nabla\phi,\vartheta)$ is given as follows
$F(\phi,\nabla\phi,\vartheta):=\frac{\gamma\vartheta}{2}|\nabla\phi|^{2}+f(\phi,\vartheta),$ (4)
The relationship between the chemical potential $\frac{\mu}{\vartheta}$, internal energy density $e$ and entropy density $s$ is then expressed via
$\frac{\mu}{\vartheta}$ $:=\delta_{\phi}\frac{F(\phi,\nabla\phi,\vartheta)}{\vartheta}=-\gamma\Delta\phi+\partial_{\phi}\tfrac{f(\phi,\vartheta)}{\vartheta}$ (5)
$s(\phi,\nabla\phi,\vartheta)$ $:=-\partial_{\vartheta}F(\phi,\nabla\phi,\vartheta)=-\frac{\gamma}{2}|\nabla\phi|^{2}-\partial_{\vartheta}f(\phi,\vartheta),$ (6)
$e(\phi,\vartheta)$ $:=F(\phi,\nabla\phi,\vartheta)+\vartheta s(\phi,\nabla\phi,\vartheta)=f(\phi,\vartheta)-\vartheta\partial_{\vartheta}f(\phi,\vartheta).$ (7)
This structure guarantees the following thermodynamic properties:
@@
$\int_{\Omega}s(\phi(t),\nabla\phi(t),\vartheta(t))=\int_{\Omega}s(\phi(0),\nabla\phi(0),\vartheta(0))+\int_{0}^{t}\mathcal{D}\left(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta}\right)$
where the entropy production $\mathcal{D}$ is given by
@@
$\mathcal{D}\left(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta}\right)=\int_{\Omega}\begin{pmatrix}\nabla\frac{1}{\vartheta}&\nabla\frac{\mu}{\vartheta}\end{pmatrix}\begin{pmatrix}\mathbf{K}&-\mathbf{C}\\ -\mathbf{C}&\mathbf{M}\end{pmatrix}\begin{pmatrix}\nabla\frac{1}{\vartheta}\\ \nabla\frac{\mu}{\vartheta}\end{pmatrix}$.
If the Onsager matrix
$\mathbf{L}=\begin{pmatrix}\mathbf{K}&-\mathbf{C}\\ -\mathbf{C}&\mathbf{M}\end{pmatrix}$, (8)
which depends on $\phi,\nabla\phi$ and $\vartheta$, i.e. $\mathbf{L}=\mathbf{L}(\phi,\nabla\phi,\vartheta)$, is positive (semi)-definite, then the entropy $s(\phi,\nabla\phi,\vartheta)$ is nondecreasing in time.
Various extensions of non-isothermal phase-field models have been developed in the literature, encompassing both conserved and non-conserved dynamics. General frameworks can be found in [32, 24]. The non-conserved case was notably addressed by Penrose and Fife [33], while the conserved formulation of Alt and Pawłow [2] uses entropy as the driving potential. These models can also be derived using the GENERIC or SNET formalism [18], or from a microforce balance perspective following Gurtin’s framework [23, 30, 29], or via variational principles such as the least action [14]. Finite-speed thermal propagation has also been addressed, e.g., in models by Caginalp [9, 13]. Further extensions to include (quasi-)incompressible fluid flow are discussed in [1, 15]. Analytical studies on these models, including the existence of weak solutions, especially in the presence of cross-coupling terms, can be found in [3, 26, 4, 12, 13].
@@
On the numerical side, an implicit Euler method using inverse temperature as a variable was first introduced in [32], and later enhanced for energy stability by González et al. [20] via the Average Vector Field (AVF) method [19]. More recently, Brunk et al. [8, 7] proposed structure-preserving schemes for such systems using the inverse temperature as the main variable. The Cahn-Hilliard and Allen-Cahn equations, known to exhibit a gradient structure, have been approximated using techniques like convex-concave splitting [38], the Scalar Auxiliary Variable (SAV) approach [37, 28], and Energy Quadratisation (EQ) [11, 42]. These methods introduce auxiliary variables to relax the energy or entropy functional. In particular, EQ methods have been applied to the internal energy equation in [22, 39].
@@
The goal of this work is to investigate the variational structure of the NCH system (1)–(3) and to develop a systematic discretisation in space and time using standard temperature variable, instead of the inverse temperature. The main contributions of this paper are:
@@
  • A reformulation of the system using the entropy equation in place of the internal energy equation to highlight the underlying variational structure.
  • A structure-preserving numerical approximation, employing conventional discretisation techniques in both space and time.
@@
The manuscript is organised as follows: In Section 2, we introduce notation and reformulate the system by replacing the internal energy equation with its entropy-based counterpart. In Section 3, we derive a variational formulation of the reformulated system suitable for standard finite element approximation. Section 4 and Section 5 present the fully discrete scheme based on a mixed explicit-implicit time integration and convex-concave splitting of the potential. Numerical results, including convergence studies and an application-motivated example, are provided in Section 6.
@@ 2 Preliminaries
Before delving into the details of our discretisation strategy and main results, we first introduce the necessary notation, state our primary assumptions, and recall some basic definitions.
@@ 2.1 Notation
We study the system (1)–(3) over a finite time interval $(0,T).$ To simplify the analysis and avoid complications related to boundary conditions, we assume a periodic spatial domain:
@@
    Let $\Omega\subset\mathbb{R}^{d}$, $d=1,2,3$ be a cube, identified with the $d$-dimensional torus $\mathcal{T}^{d}$. That is, we impose periodic boundary conditions and assume that the functions on $\Omega$ are periodic throughout the paper.
We denote by $L^{p}(\Omega)$ and $W^{k,p}(\Omega)$ the standard periodic Lebesgue and Sobolev spaces, equipped with norms $\|\cdot\|_{L^{p}}$ and $\|\cdot\|_{W^{k,p}}$, respectively. In particular, we set $H^{k}(\Omega)=W^{k,2}(\Omega)$ and use $\|\cdot\|_{H^{k}}=\|\cdot\|_{W^{k,2}}$. The duality pairing on $H^{-k}(\Omega)\times H^{k}(\Omega)$ is denoted by $\langle\cdot,\cdot\rangle$. We use the same notation for scalar product on $L^{2}(\Omega)$, which is defined as
$\langle u,v\rangle=\int_{\Omega}u\cdot v\,dx\qquad\forall u,v\in L^{2}(\Omega).$
For later use, we recall the following topological degree theorem from [16] which will be crucial in proving the existence of discrete solutions in our analysis:
@@
Let N and M be two positive integers and $C_{1},C_{2}$ and $\varepsilon$ three positive constants. We define
$V$ $=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M}\penalty 10000\ \text{such that}\penalty 10000\ \mathbf{y}>0\},$
$W$ $=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M}\penalty 10000\ \text{s.t.}\penalty 10000\ \|\mathbf{x}\|<C_{1}\penalty 10000\ \text{and}\penalty 10000\ \varepsilon<\mathbf{y}<C_{2}\}.$
Here, the notation $\mathbf{y}>c$ means that each component of $\mathbf{y}$ is greater than a constant $c\in\mathbb{R}$. Further $\|\cdot\|$ is a norm defined over $\mathbb{R}^{N}$. Let $\mathbf{b}\in\mathbb{R}^{N}\times\mathbb{R}^{M}$, $\mathbf{g}$ and $\mathbf{G}$ be two continuous functions, respectively, from $V$ and $V\times[0,1]$ to $\mathbb{R}^{N}\times\mathbb{R}^{M}$ satisfying the following conditions:
@@
    $\mathbf{G}(\cdot,1)=\mathbf{g}(\cdot)$
@@
    The topological degree of $\mathbf{G}(\cdot,0):=\mathbf{G}_{0}(\cdot)$ with respect to $W$ and $\mathbf{b}$ is equal to $d_{0}\neq 0$, i.e.
    $\mathrm{deg}(\mathbf{G}_{0},W,\mathbf{b}):=\sum_{\mathbf{x}\in\mathbf{G}_{0}^{-1}(\mathbf{b})}\mathrm{sgn}\left(\det\left(J_{\mathbf{G}_{0}}(\mathbf{x})\right)\right)=d_{0}.$
@@
    For all $\alpha\in[0,1]$, if $\mathbf{v}\in V$ is such that $\mathbf{G}(\mathbf{v},\alpha)=\mathbf{b}$ then $\mathbf{v}\in W$.
Then the topological degree of $\mathbf{G}(\cdot,1)$ with respect to $W$ and $\mathbf{b}$ is also equal to $d_{0}\neq 0$. Consequently, there exists at least one solution $\mathbf{v}\in W$, such that $\mathbf{g}(\mathbf{v})=\mathbf{b}$.
@@ 2.2 Assumptions
In this subsection, we summarize the minimal set of assumptions that will be used throughout the remainder of the manuscript.
@@
    The interface parameter $\gamma\in\mathbb{R}$ is strictly positive.
@@
    The Onsager matrix $\mathbf{L}=\mathbf{L}(\phi,\nabla\phi,\vartheta)\in\mathbb{R}^{2d\times 2d}$ from (8) is positive definite. Specifically, there exist constants $\lambda_{0},\lambda_{1}>0$ such that for all $\mathbf{\xi}\in\mathbb{R}^{2d}$
    $\lambda_{0}|\xi|^{2}\leq\mathbf{\xi}^{\top}\mathbf{L}\mathbf{\xi}\leq\lambda_{1}|\xi|^{2}.$
@@
    The driving potential $f(\cdot,\cdot):\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ is smooth. For each fixed $\phi\in\mathbb{R}$, the function $f(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is concave and diverges as $\vartheta\to 0$. Moreover, for each fixed $\vartheta>0$, the function $f(\cdot,\vartheta)$ can be decomposed into a convex and a concave part denoted by $f_{\text{vex}}$ and $f_{\text{cav}}$, such that $f(\cdot,\vartheta)=f_{\text{vex}}(\cdot,\vartheta)+f_{\text{cav}}(\cdot,\vartheta)$.
@@ 2.3 Structural identities
$\langle\partial_{t}\phi,1\rangle=\langle\partial_{t}e(\phi,\vartheta),1\rangle=0,\quad\langle\partial_{t}s(\phi,\nabla\phi,\vartheta),1\rangle=\mathcal{D}\left(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta}\right)\geq 0.$ (9)
@@
$\partial_{t}s(\phi,\nabla\phi,\vartheta)$ $=\partial_{t}\tfrac{e(\phi,\vartheta)-F(\phi,\nabla\phi,\vartheta))}{\vartheta}$
$=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{t}F(\phi,\nabla\phi,\vartheta)-\tfrac{e(\phi,\vartheta)-F(\phi,\nabla\phi,\vartheta)}{\vartheta^{2}}\partial_{t}\vartheta$
$=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{\phi}F(\phi,\nabla\phi,\vartheta)\partial_{t}\phi-\tfrac{1}{\vartheta}\partial_{\nabla\phi}F(\phi,\nabla\phi,\vartheta)\partial_{t}\nabla\phi$
$\quad-\tfrac{1}{\vartheta}\partial_{\vartheta}F(\phi,\nabla\phi,\vartheta)\partial_{t}\vartheta-\tfrac{s(\phi,\nabla\phi,\vartheta)}{\vartheta}\partial_{t}\vartheta.$
Using the relation $s(\phi,\nabla\phi,\vartheta)=-\partial_{\vartheta}F(\phi,\nabla\phi,\vartheta)$ and substituting in (4) leads to
$\partial_{t}s(\phi,\nabla\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\partial_{t}e(\phi,\vartheta)-\tfrac{1}{\vartheta}\partial_{\phi}f(\phi,\vartheta)\partial_{t}\phi-\gamma\nabla\phi\cdot\nabla\partial_{t}\phi.$
@@
$\partial_{t}s(\phi,\nabla\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})-\tfrac{\mu}{\vartheta}\partial_{t}\phi-\gamma\operatorname{div}(\partial_{t}\phi\nabla\phi).$
@@
$\partial_{t}s(\phi,\nabla\phi,\vartheta)$ $=\tfrac{1}{\vartheta}\operatorname{div}(\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\mathbf{K}\nabla\tfrac{1}{\vartheta})-\tfrac{\mu}{\vartheta}\operatorname{div}(\mathbf{M}\nabla\tfrac{\mu}{\vartheta}-\mathbf{C}\nabla\tfrac{1}{\vartheta})$
@@
$\partial_{t}s(\phi,\nabla\phi,\vartheta)$ $=\nabla\tfrac{1}{\vartheta}\cdot\mathbf{K}\nabla\tfrac{1}{\vartheta}-2\nabla\tfrac{1}{\vartheta}\cdot\mathbf{C}\nabla\tfrac{\mu}{\vartheta}+\nabla\tfrac{\mu}{\vartheta}\cdot\mathbf{M}\nabla\tfrac{\mu}{\vartheta}$
$-\operatorname{div}\left(\gamma\partial_{t}\phi\nabla\phi+\tfrac{1}{\vartheta}\mathbf{K}\nabla\tfrac{1}{\vartheta}-\tfrac{1}{\vartheta}\mathbf{C}\nabla\tfrac{\mu}{\vartheta}-\tfrac{\mu}{\vartheta}\mathbf{C}\nabla\tfrac{1}{\vartheta}+\tfrac{\mu}{\vartheta}\mathbf{M}\nabla\tfrac{\mu}{\vartheta}\right).$ (10)